Demand curve q= 10-2p

Price elasticity of demand = -2/3 at price p* how much is p*


1
Expert's answer
2021-10-14T17:06:52-0400

The price elasticity of demand at "p^*" is

"\frac{dQ_Bp^*}{dpQ_B^*} = -2 \times \frac{p^*}{10-2p^*} = \frac{-2}{3} \\

3p^* = 10-3p^* \\

5p^* = 10 \\

p^*= 2"